OpenStudy (anonymous):

Solve: 8x - y = -6 and xy = 9

OpenStudy (2bornot2b):

\[x=9/y\] Putting the above value of x in the first equation \[72/y-y=-6\implies 72-y^2=-6y\implies y^2-6y-72=0\]

OpenStudy (2bornot2b):

I hope you can solve the quadratic equation....

OpenStudy (anonymous):

So I just do the quad form and then it should give me my answer for Y?

OpenStudy (anonymous):

Awesome, thanks

OpenStudy (2bornot2b):

Yes, you will get two values of y, by solving the above quadratic equation. Now put each of them on any one of your equations, and you will get two values of x. So that's the solution, two values of x and two values of y.

OpenStudy (2bornot2b):

You are welcome! It also must be mentioned here, that the solutions must be written in such a way such that the reader can make out from your writing which x is corresponding to which y.